a little help please..
Which TWO linear equations represent perpendicular lines? Select two choices.
y = 2x + 4
x - 3y = -9
y = -2x - 3
y = ½x - 4

Answer by richwmiller(17219)

perpendicular lines have slopes whose product =-1
put the equations in the form y=mx+b
where m is the slope
fortunately the second pair is in that form and
-2*1/2=-1 they are perpendicular

the first pair y=2x+4
and x-3y=-9
add -x to both sides
-3y=-x-9
divide both sides by -3
y=-x/-3-9/-3
y=x/3+3
m=1/3
and m=2
2*1/3 does not =-1